OpenStudy (giitter):

Which is the best estimate for the product of 3,826.301 × 175.6? A. 8,000,000 B. 800,000 C. 80,000 D. 8,000

OpenStudy (giitter):

@AloneS

OpenStudy (giitter):

@alivejeremy

OpenStudy (xxhero782xx):

B

OpenStudy (mathstudent55):

3,826.301 × 175.6? 3,826.301 is close to 4,000 175.6 is close to 200 Also, 200 = 100 * 2 4,000 * 200 = 4,000 * 100 * 2 = 400,000 * 2 = 800,000